Calculate the length of one of the sides of a right angled triangle using the Pythagorean theorem. What is the length of side B of the triangle in which side A=6cm and side C=10cm ?
C squared= A squared + B squared
I keep trying numbers but can't get any to add up equally without being a decimal & having to use every number after it.
Help please?

Answer by ewatrrr(24785)

.
B^2 = C^2 - A^2
B^2 = 100cm^2 - 36cm^2
B^2 = 64cm^2 (Tossing out the negative solution for unit measure)
B = 8cm